OpenStudy (itskels2001):

Given a function f(x), let g(x) = f(x) - 10. How do the graphs of f(x) and g(x) compare? Question 5 options: g(x) is translated 10 units down. g(x) is translated 10 units up. g(x) is translated 10 units left. g(x) is translated 10 units right.

OpenStudy (itskels2001):

@mecharv b maybe? I have no idea

OpenStudy (mecharv):

Like we have already discussed, this is just a vice versa of that it is a

OpenStudy (mathmale):

Why not try a little experiement? Graph f(x)=x^2, and then graph g(x)=x^2-10. Compare the two graphs.

OpenStudy (mecharv):

@mathmale for a similar question this was done Sargent.